Effective Yield 0.05
NPER 50
Coupon 31.25
FV 1000
what formula use and calculate 657.70
1
Expert's answer
2019-12-10T08:42:21-0500

"P = C\times\frac{1 - (1+i)^{-n}}{i} + M\times(1 + i)^{-n} = 31.25\times\frac{1 - (1+0.05)^{-50}}{0.05} + 1,000\times(1 + 0.05)^{-50} = 657.7."